OpenStudy (anonymous):

Which is equivalent to (-2 + 4i)(7 + 2i)?

OpenStudy (eyust707):

-14 +28i -4i +8i^2 = -14 +24i -8

OpenStudy (eyust707):

= -22 + 24i